# A RankNet-Inspired Surrogate-Assisted Hybrid Metaheuristic for Expensive Coverage Optimization

### v1

@@ V-B3 Comparative Analysis of Runtime
The runtime of each algorithm is also investigated in our experiments. All algorithms are implemented in Python and executed under the same environment for fair comparison. MixedEGO is excluded due to its runtime exceeding 1 day across all problem scales. As shown in Fig. 10, RI-SHM achieves the shortest runtime despite employing deep learning techniques. This efficiency is attributed to two aspects: (1) PyTorch’s training environment facilitates internal parallelization, and (2) the online update strategy in RI-SHM processes only a small number of new samples, which reduces the computational costs compared to full-data updates.
@@ V-C Ablation Study
To evaluate the contributions of each component in RI-SHM, we conduct an ablation study in this part. RI-SHM-w/oL and RI-SHM-w/oG are two variants of our algorithm that exclude the local surrogate-assisted EDA and the global RankNet-based pairwise surrogate, respectively. This highlights the importance of the global and local modules. Specifically, one instance from each problem scale is randomly selected, and each variant is run by 10 times independently. The statistical results are summarized in Table VI.
Table VI shows that RI-SHM significantly outperforms its variants, particularly RI-SHM-w/oG. Leveraging a small number of samples, the global surrogate enables fast online updates to explore the solution space and identify superior solutions. Fig. 11 presents the convergence curves for two instances of medium and large scales, further validating the global surrogate’s effectiveness. The surrogate-assisted local EDA plays a complementary role by helping the algorithm escape from local optima and refine existing solutions. Although its convergence speed is relatively slow, it is adequate for the demands of the local optimization phase. Without this component, it would be challenging to achieve better coverage for sensor networks. Both the global and local modules are essential to RI-SHM’s performance, as removing either leads to a notable decline in effectiveness.
@@ V-D Impact Analysis of RankNet-based Pairwise Surrogate
To investigate the performance of the RankNet-based pairwise surrogate, we replace RI-SHM’s global surrogate with RankSVM [^5] https://gist.github.com/agramfort/2071994[44], Random Forest (RF), and RBFN [7] for comparative analysis. The first two are pairwise-based, while the latter is regression-based. To assess the influence of the global surrogate, the surrogate-assisted local EDA was removed. We follow the method described in [42] to construct training and test datasets. The initial archive is used as training data to ensure uniform training conditions across surrogate models, while offspring solutions generated by the population are used as the test set to evaluate the model’s performance in predicting the ranking between pairs of individuals. Fitness performance and prediction accuracy are used as evaluation metrics. For all the compared learning algorithms, no hyperparameter optimization is conducted for fair comparison.
Table VII presents the statistical performance and prediction accuracy across six instances of varying scales, where ”RankNet” refers to the RankNet-based pairwise global surrogate. We also conduct the Wilcoxon rank-sum test similar to Section V-B1 on the results and calculate the average rankings with the Friedman test. The analysis reveals that RankNet achieves better rankings compared to other algorithms. It attains the best statistical results and outperforms RankSVM and RBFN in accuracy. RF delivers comparable predictive accuracy for individual rankings and even outperforms RankNet in certain cases, which is due to its ensemble learning mechanism. However, RF-based optimization shows poor fitness performance, as it can identify superior individuals but struggles to capture significant changes, often leading to stagnation in local optima. RBFN performs the worst overall, highlighting the advantages of pairwise-based surrogates over regression-based ones.
@@ V-E Parameter Sensitivity Analysis
The threshold $\delta$ serves as a core component of RI-SHM, which controls the switching between global and local optimization phases to balance exploration and exploitation. At the end of each iteration, the algorithm evaluates the fitness diversity of the current population. If the fitness diversity falls below $\delta$, the algorithm switches to the alternate optimizer to detect fresh genotype. To determine the appropriate value for $\delta$, we randomly select a medium-scale instance as the test scenario. The medium-scale instance represents a balanced compromise between small and large-scale cases, making it a suitable choice for evaluating performance across varying problem scales.
@@ VI Conclusion and future work
In this paper, We reformulate coverage optimization via the lens of EOPs to create a high-fidelity realistic model, and propose a novel RI-SHM to address this EMVOP. RI-SHM is a SAEA that integrates three components. First, the RankNet-based pairwise global surrogate predicts the rankings between pairs of individuals, which enables the identification of promising individuals for real FEs. Second, the surrogate-assisted local EDA is employed to continuously refine solutions and help escape from local optima. Finally, the fitness diversity-driven strategy adaptively switches between global and local optimizers during the evolution.
To validate the effectiveness of our RI-SHM, we conduct experiments on 15 test instances across different problem scales, with dimensions ranging from 75 to 300. Comparative results against other algorithms demonstrate the superiority of RI-SHM in terms of both solution quality and robustness. Additionally, we further investigate the algorithm’s components and find our that the predictive accuracy of RankNet-based surrogate is usually superior to other models. These findings underscore the capability of RI-SHM in addressing high-dimensional EMVOPs effectively. In the future, we aim to extend RI-SHM to standard benchmark Black-Box Optimization problems and address a broader range of typical EMVOPs (i.e., NAS). Additionally, exploring coverage optimization under uncertainty is another promising direction for future research [53].
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v2

@@ V-B3 Comparative Analysis of Runtime
Our experiments also investigated the runtime of each algorithm. All algorithms are implemented in Python and executed under the same environment for fair comparison. MixedEGO is excluded due to its runtime exceeding 1 day across all problem scales. As shown in Fig. 10, RI-SHM achieves the shortest runtime despite employing deep learning techniques. This efficiency is attributed to two aspects: (1) PyTorch’s training environment facilitates internal parallelization, and (2) the online update strategy in RI-SHM processes only a small number of new samples, reducing computational costs compared to full-data updates.
@@ V-C Ablation Study
To evaluate the contributions of each component in RI-SHM, we conducted an ablation study on this part. RI-SHM-w/oL and RI-SHM-w/oG are variants of our algorithm that exclude the local surrogate-assisted EDA and the global RankNet-based pairwise surrogate, respectively. This highlights the importance of the global and local modules. Specifically, one instance from each problem scale is randomly selected, and each variant is run 10 times independently. The statistical results are summarized in Table VI.
Table VI shows that RI-SHM significantly outperforms its variants, particularly RI-SHM-w/oG. Leveraging a small number of samples, the global surrogate enables fast online updates to explore the solution space and identify superior solutions. Fig. 11 presents the convergence curves for two instances of medium and large scales, further validating the global surrogate’s effectiveness. The surrogate-assisted local EDA is complementary by helping the algorithm escape from local optima and refine existing solutions. Although its convergence speed is relatively slow, it is adequate for the demands of the local optimization phase. Without this component, achieving better coverage for sensor networks would be challenging. The global and local modules are essential to RI-SHM’s performance, as removing either leads to a notable decline in effectiveness.
@@ V-D Impact Analysis of RankNet-based Pairwise Surrogate
To investigate the performance of the RankNet-based pairwise surrogate, we replace RI-SHM’s global surrogate with RankSVM [^5] https://gist.github.com/agramfort/2071994[30], Random Forest (RF), and RBFN [11] for comparative analysis. The first two are pairwise-based, while the latter is regression-based. The surrogate-assisted local EDA was removed to assess the global surrogate’s influence. We follow the method described in [27] to construct training and test datasets. The initial archive is used as training data to ensure uniform training conditions across surrogate models. In contrast, offspring solutions generated by the population are used as the test set to evaluate the model’s performance in predicting the ranking between pairs of individuals. Fitness performance and prediction accuracy are used as evaluation metrics. For all the compared learning algorithms, no hyperparameter optimization is conducted for fair comparison.
Table VII presents the statistical performance and prediction accuracy across six instances with varying scales, where ”RankNet” refers to the RankNet-based pairwise global surrogate. We also conduct the Wilcoxon rank-sum test similar to Section V-B1 on the results and calculate the average rankings with the Friedman test. The analysis reveals that RankNet achieves better rankings than other algorithms. It attains the best statistical results and outperforms RankSVM and RBFN in accuracy. RF delivers comparable predictive accuracy for individual rankings and even outperforms RankNet in some instances due to its ensemble learning mechanism. However, RF-based optimization shows poor fitness performance, as it can identify superior individuals but struggles to capture significant changes, often leading to stagnation in local optima. RBFN performs the worst overall, highlighting the advantages of pairwise-based surrogates over regression-based ones.
@@ V-E Parameter Sensitivity Analysis
The threshold $\delta$ is a core component of RI-SHM, which controls the switching between global and local optimization phases to balance exploration and exploitation. At the end of each iteration, the algorithm evaluates the fitness diversity of the current population. If the fitness diversity falls below $\delta$, the algorithm switches to the alternate optimizer to detect fresh genotypes. To determine the appropriate value for $\delta$, we randomly select a medium-scale instance as the test scenario. The medium-scale instance represents a balanced compromise between small and large-scale cases, making it suitable for evaluating performance across varying problem scales.
@@ VI Conclusion and future work
In this paper, We reformulate coverage optimization via the lens of EOPs to create a high-fidelity realistic model, and propose a novel RI-SHM to address this EMVOP. It generally integrates three components. First, the RankNet-based pairwise global surrogate predicts the rankings between pairs of individuals, which enables the identification of promising individuals for real FEs. Second, the surrogate-assisted local EDA is employed to continuously refine solutions and help escape from local optima. Finally, the fitness diversity-driven strategy adaptively switches between global and local optimizers during the evolution.
To validate the effectiveness of RI-SHM, we conduct experiments on 15 test instances across different problem scales, with dimensions ranging from 75 to 300. Comparative results against other algorithms demonstrate the superiority of RI-SHM in terms of both solution quality and robustness. Additionally, we further investigate the algorithm’s components and find that the predictive accuracy of RankNet-based surrogate is usually superior to other models. These findings underscore the capability of RI-SHM to address high-dimensional EMVOPs effectively. In the future, we aim to extend RI-SHM to standard benchmark Black-Box Optimization problems and address a broader range of typical EMVOPs (i.e., NAS). Additionally, exploring coverage optimization under uncertainty is another promising direction [38].